# Diffusion Beats Autoregressive in Data-Constrained Settings

### v6

@@ 1 Introduction
Training large language models (LLMs) on massive corpora of internet text has become the driver of recent AI breakthroughs (brown2020language, ; radford2019language, ; touvron2023llama, ). This progress has been fueled by scaling two core resources proportionately: compute and data  (kaplan2020scaling, ; hoffmann2022training, ). While compute availability has steadily grown—enabled by advances in hardware and the construction of larger data centers—the growth in high-quality data has been comparatively stagnant. Recent projections, such as those by Villalobos et.al. (villalobos2022will, ), estimate that the global supply of publicly available, human-generated data may be exhausted by 2028, posing a serious bottleneck to further scaling. This looming constraint makes it increasingly important to develop modeling strategies that are more data-efficient. Furthermore, there are several domains, such as robotics and healthcare, where the data, not compute, is a scarce resource even to begin with.
LLM development has so far been dominated by autoregressive (AR) models, which factorize the joint distribution of text in a fixed left-to-right order. While this modeling approach has delivered state-of-the-art performance across a range of benchmarks, it remains unclear whether it is the optimal strategy going forward. Recently, diffusion-based models—specifically masked diffusion models  austin2021structured ; sahoo2024simple ; lou2310discrete ; shi2024simplified ; arriola2025block —have emerged as an alternative strategy, framing text generation as an iterative masked denoising process rather than next-token prediction. At each step, the model predicts a randomly masked subset of tokens conditioned on the remaining ones, implicitly averaging over many conditional prediction orders instead of committing to one. Although these models have demonstrated similar scaling behavior to AR models (nie2024scaling, ; swerdlow2025unidisc, ), their practical benefits have, so far, been modest—largely due to their high training compute requirements.
@@
But a critical nuance is often overlooked: these comparisons are based entirely on single-epoch training, where each token is seen only once. This conflates compute efficiency with data efficiency, making it unclear whether diffusion models truly need 16× more compute—or simply 16× more data.
@@
    Exposure to different token orderings helps explain diffusion’s data efficiency. By adding explicit data augmentations to AR training, we find that diffusion models’ advantage arises from their exposure to a diverse set of token orderings. Essentially, the randomized masking in diffusion’s objective serves as implicit data augmentation, allowing it to generalize beyond the fixed left-to-right factorization of AR models. (Section 4.5)
@@
2 Related Work
Deep Learning in Data-Constrainted Settings.
Deep learning progress has been largely driven by the scaling of both data and compute. However, recent analyses suggest we may soon face a data bottleneck that could inhibit continued advancement (villalobos2022will, ). In language modeling, the dominant paradigm has been autoregressive (AR) models (vaswani2017attention, ; radford2019language, ; brown2020language, ), which are typically trained for a single epoch to maximize exposure to unique tokens (hoffmann2022training, ). In light of looming data constraints, Muennighoff et al.(muennighoff2023scaling, ) show that AR models can still benefit from data reuse: training for up to four epochs on repeated data achieves performance nearly on par with training on fresh data, suggesting an effective strategy for improving data efficiency. In contrast, computer vision has long embraced multi-epoch training along with aggressive data augmentation—such as random cropping, flipping, and color jittering—to expand effective dataset size and improve generalization(shorten2019survey, ; yang2022image, ), particularly for discriminative tasks like classification and detection. Despite these practices, data efficiency in generative modeling remains underexplored, and the trade-offs between leading paradigms such as diffusion and AR models under constrained data regimes are still poorly understood.
Diffusion-Based Language Models.
[…]
@@ 3 Method
Our objective is to determine whether masked diffusion language models are more effective than standard autoregressive models in data-constrainted settings. The main difference between AR and diffusion models is the way they factorize the joint distribution of the sequence. Masked diffusion factorizes the joint distribution of the sequence in a random order, while AR factorizes the joint distribution of the sequence in a left-to-right order. To isolate the impact of this, we keep the architecture and data pipeline fixed across both families and vary only the factorization of the joint distribution.
@@ 3.1 Preliminaries:
In Autoregressive LLMs (touvron2023llama, ; radford2019language, ; brown2020language, ) we predict each token based on a growing prefix of prior tokens, defining a left-to-right factorization of the sequence probability:
@@
Masked diffusion language models  (austin2021structured, ; sahoo2024simple, ; nie2024scaling, ; swerdlow2025unidisc, ) treat generation as iterative denoising. For each training sequence $x=(x_{1},\dots,x_{L})$ we
@@
$p_{\text{Diffusion}}(x\mid\tilde{x})\;=\;\prod_{i\in\mathcal{M}}p_{\theta}\!\bigl{(}x_{i}\,\mid\,\tilde{x}\bigr{)}$.
Because the mask pattern is resampled for every example, the model is implicitly trained on a vast collection of token–ordering tasks; the standard left-to-right ordering used by AR models is just one ordering within this ensemble. The absence of a causal mask allows each prediction to attend to both past and future unmasked tokens, making the factorization fundamentally non-sequential.
@@ 3.2 Modeling Details for AR and Masked Diffusion
Our goal is to isolate the impact of the factorization—fixed left-to-right versus random-order denoising—while keeping every other design choice constant. Unless noted otherwise, both model families share the same Transformer backbone (GPT-2 style with rotary positional embeddings, RoPE (su2021roformer, )) and identical training hyper-parameters across the full parameter sweep (7 M – 2 B).
@@
$\mathcal{L}_{\text{AR}}\;=\;-\sum_{j=2}^{L}\log p_{\theta}\!\bigl{(}x_{j}\mid x_{<j}\bigr{)}.$
@@
$\mathcal{L}_{\text{Diffusion}}\;=\;-\mathbb{E}_{r}\;\mathbb{E}_{\tilde{x}\sim q_{r}}\;\frac{1}{r}\sum_{i\in\mathcal{M}}\log p_{\theta}\!\bigl{(}x_{i}\mid\tilde{x}\bigr{)}$.
Beyond the attention mechanism and input corruption, all other variables are held constant. We follow the hyperparameter configuration proposed by Muennighoff et al. muennighoff2023scaling for all training runs. In particular, we use a dynamic learning rate schedule that adapts to the number of training epochs. The only distinctions between AR and diffusion models in our implementation are:
  1. 1.
    Attention mechanism: Causal attention for AR; full self-attention for masked diffusion.
  2. 2.
    Prediction target: AR models predict the next token; diffusion models predict the masked tokens.
@@ 3.3 Scaling Framework in Data-Constrained Settings
Classical scaling laws, such as those proposed by (kaplan2020scaling, ; hoffmann2022chinchilla, ), model validation loss as a function of total parameters ($N$) and training tokens ($D$), assuming all data is unique. These laws have been instrumental in guiding compute-optimal training of language models. However, this assumption becomes unrealistic as the community approaches the limits of high-quality text data available on the internet.
To address this, Muennighoff et al.muennighoff2023scaling extend the Chinchilla framework to explicitly account for repeated data — a common necessity in data-constrained regimes. They show that repeating training data beyond a few epochs yields diminishing returns and propose a new scaling law that incorporates the decaying utility of repeated tokens.
We briefly outline their formulation below.
Definitions:
  • $U$: number of unique tokens available for training,
  • $E$: number of epochs (i.e., how many times each unique token is reused),
  • $D=U\cdot E$: total number of tokens seen by the model.
@@
A symmetric formulation is applied to model parameters for mathetmatical convenience which is used to define $N^{\prime}$. Finally, a modified Chinchilla-style loss function incorporates these effective quantities $N^{\prime}$ and $D^{\prime}$:
@@ 3.4 Training setup
We use the English C4 corpus (raffel2020exploring, ), tokenized with the GPT-2 BPE vocabulary and truncated or padded to 2048 tokens per sequence. We consider unique-token budgets of $U\in\{25,50,100\}$M and train for up to 800 epochs (80B tokens total). Models are trained ranging from 7M to 2.5B parameters, following the Chinchilla scaling strategy where both width and depth are increased proportionally. The detailed architectural configurations of each model are provided in Appendix 10. For all training runs, we adopt the hyperparameter configuration introduced by Muennighoff et al. muennighoff2023scaling . This may provide a slight advantage to autoregressive models, as these hyperparameters were originally tuned for that family. For all models, we use the following hyperparameters: batch size of 256 sequences, AdamW optimizer with $\beta_{1}{=}0.9$, $\beta_{2}{=}0.95$, $\epsilon{=}10^{-8}$, a learning rate schedule with peak 2e-4, minimum 2e-5, 1% warm-up, cosine decay, weight decay 0.1, and gradient clipping of 1.0.
@@ 4.1 Does Diffusion Beat AR in Data-Constrained Settings?
In other words: is diffusion limited by compute efficiency or by data efficiency?
@@
We train a large suite of AR and diffusion models across three unique data regimes—25M, 50M, and 100M tokens—and a wide range of training compute budgets. In Figure 1, we report empirical validation loss as a function of training FLOPs for the 50M and 100M regimes; results for the 25M setting are shown in Appendix Figure 9. We find that AR models initially outperform diffusion models when trained with the compute-optimal budget prescribed by Chinchilla scaling laws (denoted by the solid vertical line). However, this advantage disappears as training continues beyond this point. When models are allowed to train for additional epochs on repeated data, diffusion models consistently surpass AR models in validation loss across all data regimes. These findings indicate that the previously observed inefficiency of diffusion models is largely a consequence of evaluating them solely in the single-epoch regime. In data-constrained settings with repeated exposures, diffusion models extract significantly more value from the same data than their AR counterparts.
@@
To contextualize these results, we highlight two key configurations in Figure 2 for each model family: the compute-optimal point for single-epoch training, as identified by prior scaling law analyses hoffmann2022chinchilla ; nie2019adversarial (marked with a colored star in the bottom-left), and the best validation loss achieved under extended multi-epoch training (marked with a black star). At the compute-optimal point, which corresponds to training for a single epoch, diffusion models perform substantially worse than autoregressive models (10.65 vs. 7.07), consistent with prior findings that diffusion performs worse initially. However, as training is extended to hundreds of epochs, diffusion models continue to improve and eventually achieve a lower validation loss (3.55) than the best AR models (3.71). While AR models begin to overfit as training progresses, diffusion models show no signs of overfitting within our budget.
@@
Figure 2: Validation loss contours over epochs and model sizes for autoregressive (left) and diffusion (right) models, trained on 100M unique tokens. Each plot shows validation loss as a function of training epochs (x-axis) and model parameters (y-axis). The colored star marks the compute-optimal point for single-epoch training, as predicted by prior scaling laws hoffmann2022chinchilla ; nie2019adversarial , and the black star indicates the lowest validation loss achieved through extended multi-epoch training. In the single-epoch regime, diffusion models perform worse than AR models (10.65 vs. 7.07). However, when trained longer, diffusion models achieve a substantially lower final loss (3.55 vs. 3.71). This corresponds to a 67% reduction in loss for diffusion models compared to just 48% for AR models, highlighting their superior ability to leverage repeated data. These results underscore that diffusion models require significantly more training—both in epochs and compute—to realize their advantages in data-constrained settings.
@@ 4.2 Fitting Data-Constrained Scaling Laws
Beyond the overall fit, we extract two key parameters from the scaling laws: $R_{D}^{*}$, which characterizes the effective half-life of data reuse—i.e., the number of epochs after which additional training on repeated data yields diminishing returns—and $R_{N}^{*}$, which indicates the optimal model size for a given data budget. Our results reveal a sharp contrast in data reuse half-lives: diffusion models exhibit an $R_{D}^{*}$ of 512.85, compared to just 31.93 for autoregressive models. A higher $R_{D}^{*}$ implies that a model can benefit from many more repeated exposures before saturating. This suggests that diffusion models continue to improve across hundreds of epochs, while AR models quickly saturate—highlighting the superior data efficiency of diffusion models in data-constrained regimes.
@@
(a) AR Training Curves.
(b) Diffusion Training Curves.
Figure 5: Predicted validation loss for AR models (left) and Diffusion models (right) under compute-optimal settings, extrapolated to larger compute budgets. Dotted lines indicate the hypothetical case where repeated data is as valuable as new data. For AR, this holds up to about 4 epochs; for diffusion, up to 100 epochs, showing that diffusion models are much more robust to data repetition.
@@ 4.3 When to Use Diffusion over AR?
where $C$ is total training compute and $U$ is the number of unique tokens. Positive values favor AR; negative values favor diffusion. The critical compute $C_{\mathrm{crit}}(U)$ is the point where the models perform equally:
$\Delta\mathcal{L}(C_{\mathrm{crit}},U)=0.$
@@
Figure 6(b) plots the critical compute frontier $C_{\mathrm{crit}}(U)$—the compute required for diffusion to match AR at a given unique token count $U$. This frontier follows a power law:
$C_{\mathrm{crit}}(U)\propto U^{2.174}.$
The linear fit in log-log space is:
@@
The green dashed line shows the fitted curve, and the blue crosses represent empirical crossover points—where diffusion matches AR performance in experiments. These points align closely with the predicted frontier, confirming our fitted equation’s accuracy.
@@ 4.4 Downstream Results
We evaluate the best-performing diffusion and autoregressive (AR) models on several downstream benchmarks to assess whether the gains in validation loss translate to practical improvements in generalization.
Motivated by the critical compute threshold equation identified in Section 4.3, we scale the training data to 500M unique tokens and train a 2.3B parameter diffusion model using the compute budget predicted by the critical compute limit. We train the model for 130 epochs, during which we observe no signs of convergence. We terminate the training due to compute constraints.
Across a diverse set of tasks and data scales, diffusion models consistently outperform their AR counterparts. This validates our findings in Section 4.3, also confirms that the data efficiency gains observed in validation loss translate into stronger downstream performance. Additional results, including performance on textbook-style datasets, are provided in Table 3 in the Appendix.
Table 2: Downstream Results for the best autoregressive and diffusion trained in different data-constrainted settings. We report the results for the models with the best validation loss in 100M and 500M unique data regime.To better understand the difficulty of each benchmark we also report the accuracy of random baseline
Benchmarks Random Baseline 100M unique tokens 500M unique tokens
AR Diffusion AR Diffusion
ARC-Easy clark2018think 25.00 35.63 37.84 43.79 45.95
BoolQ clark2019boolq 50.00 46.00 49.38 51.87 55.26
COPA roemmele2011choice 50.00 56.33 59.00 67.00 64.83
HellaSwag zellers2019hellaswag 25.00 27.37 30.24 32.28 35.33
PiQA 50.00 60.94 60.72 65.71 65.61
RACE lai2017race 25.00 25.28 28.96 28.28 31.44
WinoGrande XL sakaguchi2021winogrande 50.00 48.87 50.97 50.61 51.51
SciQ SciQ 25.00 58.05 68.67 67.82 79.13
Lambada paperno2016lambada 00.00 10.91 15.19 15.07 22.30

### v7

@@ 1 Introduction
Training large language models (LLMs) on massive corpora of internet text has become the driver of recent AI breakthroughs (brown2020language, ; radford2019language, ; touvron2023llama, ). This progress has been fueled by scaling two core resources: compute and data  (kaplan2020scaling, ; hoffmann2022training, ). While compute availability is steadily growing—enabled by advances in hardware and the construction of larger data centers—the growth in high-quality data is stagnating. Recent projections, such as those by Villalobos et.al. (villalobos2022will, ), estimate that the global supply of publicly available, human-generated data may be exhausted in the coming years, posing a serious bottleneck to further scaling. This looming constraint makes it increasingly important to develop modeling strategies that are more sample-efficient. Furthermore, there are several domains, such as robotics and healthcare, where the data, not compute, is a scarce resource even to begin with.
LLM development has so far been dominated by autoregressive (AR) models, which factorize the joint distribution of text in a fixed left-to-right order. While this modeling approach has delivered state-of-the-art performance across a range of benchmarks, it remains unclear whether it is the optimal strategy going forward. Recently, diffusion-based models—specifically masked diffusion models  austin2021structured ; shi2024simplified ; lou2310discrete ; swerdlow2025unidisc ; arriola2025block —have emerged as an alternative strategy, framing text generation as an iterative masked denoising process rather than next-token prediction. At each step, the model predicts a randomly masked subset of tokens conditioned on the remaining ones, implicitly averaging over many conditional prediction orders instead of committing to one. Although these models have demonstrated similar scaling behavior to AR models (nie2024scaling, ; swerdlow2025unidisc, ), their practical benefits have, so far, been modest—largely due to their high training compute requirements.
@@
But a critical nuance is often overlooked: these comparisons are based entirely on single-epoch training, where each token is seen only once. This conflates compute efficiency with sample efficiency, making it unclear whether diffusion models truly need 16× more compute—or simply 16× more data.
@@
    Exposure to different token orderings helps explain diffusion’s sample efficiency. By adding explicit data augmentations to AR training, we find that diffusion models’ advantage arises from their exposure to a diverse set of token orderings. Essentially, the randomized masking in diffusion’s objective serves as implicit data augmentation, allowing it to generalize beyond the fixed left-to-right factorization of AR models. (Section 3.5)
@@ 2 Method
Our objective is to determine whether masked diffusion language models are more effective than standard autoregressive models in data-constrainted settings. For studying this, we keep the core architecture and data pipeline fixed across both families.
@@ 2.1 Preliminaries:
In Autoregressive LLMs (touvron2023llama, ; radford2019language, ; brown2020language, ) each token is predicted based on a growing prefix of prior tokens, defining a left-to-right factorization of the sequence probability:
@@
Masked diffusion language models  (austin2021structured, ; shi2024simplified, ; nie2024scaling, ; swerdlow2025unidisc, ) treat generation as iterative denoising. For each training sequence $x=(x_{1},\dots,x_{L})$ diffusion models
@@
[…]
Because the mask pattern is resampled for every example, the model is implicitly trained on a vast collection of token–ordering tasks. The absence of a causal mask allows each prediction to attend to both past and future unmasked tokens.
@@ 2.2 Modeling Details for AR and Masked Diffusion
Both model families share the same Transformer backbone (GPT-2 style with rotary positional embeddings, RoPE (su2021roformer, )).
@@
$\mathcal{L}_{\text{AR}}\;=\;-\sum_{j=1}^{L}\log p_{\theta}\!\bigl(x_{j}\mid x_{<j}\bigr).$
@@
[…]
which can be interpreted as an evidence lower bound (ELBO) on the data log-likelihood, and thus the loss provides an upper bound on the true negative log-likelihood.
Beyond the attention mechanism and input corruption, all other variables are held constant. We follow the hyperparameter configuration proposed by Muennighoff et al. muennighoff2023scaling for all training runs. In particular, we use a dynamic learning rate schedule that adapts to the number of training epochs. For more detailed information on both model families please refer to related work Section in Appendix 6
@@ 2.3 Scaling Framework in Data-Constrained Settings
Classical scaling laws, such as those proposed by (kaplan2020scaling, ; hoffmann2022chinchilla, ), model validation loss as a function of total parameters ($N$) and training tokens ($D$), assuming all data is unique that is single epoch regime.
Muennighoff et al.muennighoff2023scaling extend the Chinchilla framework to explicitly account for repeated data — a common necessity in data-constrained regimes. They show that repeating training data beyond a few epochs yields diminishing returns and propose a new scaling law that incorporates the decaying utility of repeated tokens.
We briefly outline their formulation below. Let $U$ denote the number of unique tokens, $E$ the number of epochs (how many times each unique token is reused), and $D=U\cdot E$ the total number of tokens seen during training.
@@
A symmetric formulation is applied to model parameters for mathematical convenience which is used to define $N^{\prime}$. Finally, a modified Chinchilla-style loss function incorporates these effective quantities $N^{\prime}$ and $D^{\prime}$:
@@ 2.4 Training setup
We use the English C4 corpus (raffel2020exploring, ), tokenized with the GPT-2 BPE vocabulary and truncated or padded to 2048 tokens per sequence. We consider unique-token budgets of $U\in\{25,50,100\}$M and train for up to 800 epochs (80B tokens total). Models are trained ranging from 7M to 2.5B parameters, following the Chinchilla scaling strategy where both width and depth are increased proportionally. The detailed architectural configurations of each model are provided in Appendix 9. For all training runs, we adopt the hyperparameter configuration introduced by Muennighoff et al. muennighoff2023scaling . This may provide a slight advantage to autoregressive models, as these hyperparameters were originally tuned for that family. For details on hyperparameters please refer to Section 8 in Appendix.
@@ 3.1 Does Diffusion Beat AR in Data-Constrained Settings?
In other words: is diffusion limited by compute efficiency or by sample efficiency?
@@
In Figure 1, we report empirical validation loss as a function of training FLOPs for the 50M and 100M regimes; results for the 25M setting are shown in Appendix Figure 9. We find that AR models initially outperform diffusion models when trained with the compute-optimal budget prescribed by Chinchilla scaling laws (denoted by the solid vertical line). However, this advantage disappears as training continues beyond this point. When models are allowed to train for additional epochs on repeated data, diffusion models consistently surpass AR models in validation loss across all data regimes. These findings indicate that the previously observed inefficiency of diffusion models is largely a consequence of evaluating them solely in the single-epoch regime. In data-constrained settings with repeated exposures, diffusion models extract significantly more value from the same data than their AR counterparts.
@@
To contextualize these results, we highlight two key configurations in Figure 2 for each model family: the compute-optimal point for single-epoch training, as identified by prior scaling law analyses hoffmann2022chinchilla ; nie2019adversarial (marked with a colored star in the bottom-left), and the best validation loss achieved under extended multi-epoch training (marked with a black star). At the compute-optimal point, which corresponds to training for a single epoch, diffusion models perform substantially worse than autoregressive models (10.65 vs. 7.07), consistent with prior findings that diffusion performs worse initially. However, as training is extended to hundreds of epochs, diffusion models continue to improve and eventually achieve a lower validation loss (3.55) than the best AR models (3.71).
@@
Figure 2: Validation loss contours over epochs and model sizes for autoregressive (left) and diffusion (right) models, trained on 100M unique tokens. Each plot shows validation loss as a function of training epochs (x-axis) and model parameters (y-axis). The colored star marks the compute-optimal point for single-epoch training, as predicted by prior scaling laws hoffmann2022chinchilla ; nie2019adversarial , and the black star indicates the lowest validation loss achieved through extended multi-epoch training. In the single-epoch regime, diffusion models perform worse than AR models (10.65 vs. 7.07). However, when trained longer, diffusion models achieve a substantially lower final loss (3.55 vs. 3.71). This corresponds to a 67% reduction in loss for diffusion models compared to just 48% for AR models, highlighting their superior ability to leverage repeated data.
@@ 3.2 Fitting Data-Constrained Scaling Laws
(a) AR Training Curves.
(b) Diffusion Training Curves.
Figure 5: Predicted validation loss for AR (left) and Diffusion models (right) under compute-optimal settings, extrapolated to larger compute budgets. Dotted lines show the hypothetical case where repeated data equals new data. For AR, this holds up to $\approx$4 epochs; for diffusion, up to $\approx$100 epochs—showing diffusion’s greater robustness to data repetition. Note that loss values between AR and diffusion are not directly comparable, as they’re extrapolated from scaling laws with different data-entropy terms ($E_{0}$). In Section 3.5, we ignore this factor during comparison.
@@
Beyond the overall fit, we extract two key parameters from the scaling laws: $R_{D}^{*}$, which characterizes the effective half-life of data reuse—i.e., the number of epochs after which additional training on repeated data yields diminishing returns—and $R_{N}^{*}$, which indicates the optimal model size for a given data budget. Our results reveal a sharp contrast in data reuse half-lives: diffusion models exhibit an $R_{D}^{*}$ of 512.85, compared to just 31.93 for autoregressive models. A higher $R_{D}^{*}$ implies that a model can benefit from many more repeated exposures before saturating. This suggests that diffusion models continue to improve across hundreds of epochs, while AR models quickly saturate—highlighting the superior sample efficiency of diffusion models in data-constrained regimes.
@@ 3.3 When to Use Diffusion over AR?
where $C$ is total training compute and $U$ is the number of unique tokens. Positive values favor AR; negative values favor diffusion. The critical compute $C_{\mathrm{crit}}(U)$ is the point where the models perform equally: $\Delta\mathcal{L}(C_{\mathrm{crit}},U)=0.$
@@
Figure 6(b) plots the critical compute frontier $C_{\mathrm{crit}}(U)$—the compute required for diffusion to match AR at a given unique token count $U$. This frontier follows a power law: $C_{\mathrm{crit}}(U)\propto U^{2.174}.$ The linear fit in log-log space is:
@@
The dark green dashed line shows the fitted curve, and the blue stars represent empirical crossover points—where diffusion matches AR performance in experiments. These points align closely with the predicted frontier, confirming our fitted equation’s accuracy.
@@ 3.4 Downstream Results
Benchmarks Random Baseline AR AR (Flop matched) Diffusion
ARC-Easy clark2018think 25.00 35.63 35.35 37.84
BoolQ clark2019boolq 50.00 46.00 38.23 49.38
COPA roemmele2011choice 50.00 56.33 54.00 59.00
HellaSwag zellers2019hellaswag 25.00 27.37 29.03 30.24
PiQA 50.00 60.94 61.64 60.72
RACE lai2017race 25.00 25.28 24.88 28.96
WinoGrande XL sakaguchi2021winogrande 50.00 48.87 49.41 50.97
SciQ SciQ 25.00 58.05 50.50 68.67
Lambada paperno2016lambada 00.00 10.91 5.53 15.19
@@
Table 2: Downstream results for the best-performing (as per validation loss) and flop-matched autoregressive (AR) and diffusion models trained with 100M unique tokens. Random baselines are reported for reference.
We evaluate the best-performing diffusion and autoregressive (AR) models, selected based on their validation loss, across several downstream benchmarks to examine whether lower validation loss translates to improved generalization.
Since AR models tend to overfit much earlier, we additionally evaluate flop-matched overfitted AR models trained for the same number of epochs as their diffusion counterparts. On a few set of benchmarks, these overfitted AR models slightly outperform their best-validation variants; however, across almost all tasks, diffusion models consistently achieve the highest downstream performance.
Additional results, with 500M unique tokens and other downstream datasets, are provided in Table 3 and Table 4 in the Appendix.
Across a diverse set of tasks and data scales, diffusion models consistently outperform their AR counterparts.